# Reveal-or-Obscure: A Differentially Private Sampling Algorithm for Discrete Distributions

### v1

@@ V Conclusion
Since the denominator is always negative, the right side of (60) will be positive only when the numerator is also negative, i.e.,
$w_{m}=-\frac{1}{n}-\frac{m}{n}+\frac{m}{n}e^{\epsilon}<0\Rightarrow m<\frac{1}{e^{\epsilon}-1}.$ (61)
For $\tilde{m}=m+1$ and $\tilde{m}=m-1$, we obtain two more inequalities,
$u_{m}q_{m+1}\leq v_{m}q_{m}+w_{m},$ (62)
$u_{m}q_{m-1}\leq v_{m}q_{m}+w_{m}.$ (63)
Therefore, for Algorithm 2 to achieve $\epsilon$-DP, the function $q_{m}$ must satisfy the inequalities in (60), (62), and (63). Defining the initial value $q_{0}$ as in (9), and assuming (63) to be an equality for $m\geq 1$, we arrive at the expression
$q_{m}=\frac{u_{m}}{v_{m}}q_{m-1}-\frac{w_{m}}{v_{m}}.$ (64)
Note that, since $q_{m}$ is a probability, we must have $0\leq q_{m}\leq 1$ for all $m$. Therefore, (64) is valid only when the right side is positive. The final form of $q_{m}$ is thus
$q_{m}=\max\left\{0,\frac{u_{m}}{v_{m}}q_{m-1}-\frac{w_{m}}{v_{m}}\right\}.$ (65)
The proofs of (64) satisfying the conditions in (60) and (62) are provided in Lemma 3 and 4. The proof that the function $q_{m}$ is non-increasing is provided in Lemma 2. When $m=0$, $q_{m}$ is equal to that of the fixed $q$ of Algorithm 1. As $m$ increases, the input distribution gets closer to $\text{Unif}[1:k]$, and Algorithm 2 obscures the output distribution with smaller probability.
-D Additional Proofs
Lemma 2
For any $k\geq 2,\epsilon>0$, and datasets of size $n$ where $n>k$, the function $q_{m}$ is non-increasing.
@@
Since the function $q_{m}:\left\{0,1,\dots,\left\lfloor\frac{n}{k}\right\rfloor\right\}\mapsto[0,1]$ is discrete, we need to show that
$q_{m}\leq q_{m-1},$ (66)
for all $m\in\left\{0,1,\dots,\left\lfloor\frac{n}{k}\right\rfloor\right\}$. From Section IV, we have, for $m\geq 1$ and positive $q_{m}$,
$v_{m}q_{m}$ $=u_{m}q_{m-1}-w_{m}$ (67)
$\Rightarrow v_{m}q_{m}-v_{m}q_{m-1}$ $=u_{m}q_{m-1}-w_{m}-v_{m}q_{m-1}$ (68)
$\Rightarrow q_{m}-q_{m-1}$ $=\frac{u_{m}-v_{m}}{v_{m}}q_{m-1}-\frac{w_{m}}{v_{m}}$ (69)
$\Rightarrow q_{m}-q_{m-1}$ $=\left(\frac{u_{m}}{v_{m}}-1\right)q_{m-1}-\frac{w_{m}}{v_{m}}.$ (70)
It suffices to show that the right side of (70) is negative. Substituting (10) and (11) into the first coefficient, we have
$<0.$ (74)
Similarly, substituting (12) into the second coefficient, we have
$>0,$ (77)
for $m>\left\lfloor\frac{1}{e^{\epsilon}-1}\right\rfloor$. Since $q_{m}$ has a lower bound 00 in this range, the right side of (70) is
$\left(\frac{u_{m}}{v_{m}}-1\right)q_{m-1}-\frac{w_{m}}{v_{m}}<0.$ (78)
It still remains to show that $q_{m}-q_{m-1}<0$ for $m=1,2,\dots,\left\lfloor\frac{1}{e^{\epsilon}-1}\right\rfloor$. Recall from (60) in Appendix -C that, in this range of $m$, $q_{m}$ has a non-zero, positive lower bound. Substituting this bound to the right side of (70), we have
$\left(\frac{u_{m}}{v_{m}}-1\right)\frac{w_{m-1}}{u_{m-1}-v_{m-1}}-\frac{w_{m}}{v_{m}}$
$=\frac{-\frac{1}{n}-\left(\frac{1}{k}-\frac{m}{n}\right)(e^{\epsilon}-1)}{e^{\epsilon}\left(\frac{1}{k}-\frac{m}{n}\right)}\frac{\frac{m-1}{n}(e^{\epsilon}-1)-\frac{1}{n}}{\left(\frac{m-1}{n}-\frac{1}{k}\right)(e^{\epsilon}-1)-\frac{1}{n}}-\frac{w_{m}}{v_{m}}$ (79)
$=\frac{-\frac{1}{n}-\left(\frac{1}{k}-\frac{m}{n}\right)(e^{\epsilon}-1)}{\left(\frac{m-1}{n}-\frac{1}{k}\right)(e^{\epsilon}-1)-\frac{1}{n}}\frac{\frac{m-1}{n}(e^{\epsilon}-1)-\frac{1}{n}}{e^{\epsilon}\left(\frac{1}{k}-\frac{m}{n}\right)}-\frac{w_{m}}{v_{m}}$ (80)
$=C\frac{-\frac{1}{n}+\frac{m}{n}(e^{\epsilon}-1)-\frac{1}{n}(e^{\epsilon}-1)}{e^{\epsilon}\left(\frac{1}{k}-\frac{m}{n}\right)}-\frac{w_{m}}{v_{m}}$ (81)
$=C\left[\frac{-\frac{1}{n}+\frac{m}{n}(e^{\epsilon}-1)}{e^{\epsilon}\left(\frac{1}{k}-\frac{m}{n}\right)}-\frac{\frac{1}{n}(e^{\epsilon}-1)}{e^{\epsilon}\left(\frac{1}{k}-\frac{m}{n}\right)}\right]-\frac{w_{m}}{v_{m}}$ (82)
$=C\left[\frac{w_{m}}{v_{m}}-\frac{e^{\epsilon}-1}{e^{\epsilon}\left(\frac{n}{k}-m\right)}\right]-\frac{w_{m}}{v_{m}}.$ (83)
Here, $\frac{w_{m}}{v_{m}}<0$ for the relevant range of $m$. Moreover, since $\frac{n}{k}>m$, $\frac{e^{\epsilon}-1}{e^{\epsilon}\left(\frac{n}{k}-m\right)}$ is positive and less than $1$. Finally, the coefficient,
$C$ $=\frac{-\frac{1}{n}-\left(\frac{1}{k}-\frac{m}{n}\right)(e^{\epsilon}-1)}{\left(\frac{m-1}{n}-\frac{1}{k}\right)(e^{\epsilon}-1)-\frac{1}{n}}$ (84)
$=\frac{-\frac{1}{n}-\left(\frac{1}{k}-\frac{m-1+1}{n}\right)(e^{\epsilon}-1)}{\left(\frac{m-1}{n}-\frac{1}{k}\right)(e^{\epsilon}-1)-\frac{1}{n}}$ (85)
$=\frac{\left(\frac{m-1}{n}-\frac{1}{k}\right)(e^{\epsilon}-1)-\frac{1}{n}+\frac{1}{n}(e^{\epsilon}-1)}{\left(\frac{m-1}{n}-\frac{1}{k}\right)(e^{\epsilon}-1)-\frac{1}{n}}$ (86)
$=1+\frac{\frac{1}{n}(e^{\epsilon}-1)}{\left(\frac{m-1}{n}-\frac{1}{k}\right)(e^{\epsilon}-1)-\frac{1}{n}}$ (87)
$=1-\frac{e^{\epsilon}-1}{1+n(e^{\epsilon}-1)\left(\frac{1}{k}-\frac{m-1}{n}\right)}$ (88)
$<1.$ (89)
Therefore, we have
$C\left[\frac{w_{m}}{v_{m}}-\frac{e^{\epsilon}-1}{e^{\epsilon}\left(\frac{n}{k}-m\right)}\right]<\frac{w_{m}}{v_{m}}$ (90)
$\Rightarrow C\left[\frac{w_{m}}{v_{m}}-\frac{e^{\epsilon}-1}{e^{\epsilon}\left(\frac{n}{k}-m\right)}\right]-\frac{w_{m}}{v_{m}}<0$ (91)
Thus, for all $m$,
$q_{m}\leq q_{m-1},$ (92)
i.e., $q_{m}$ is non-increasing. ∎
Lemma 3
For $m=1,2,\dots,\left\lfloor\frac{1}{e^{\epsilon}-1}\right\rfloor$, and $u_{m},v_{m}$, the function $q_{m}$ satisfies the inequality condition
$q_{m}>\frac{w_{m}}{u_{m}-v_{m}}.$ (93)
Proof:
Let,
$t_{m}=\frac{w_{m}}{u_{m}-v_{m}}$ $=\frac{\frac{w_{m}}{v_{m}}}{\frac{u_{m}}{v_{m}}-1}$ (94)
$\Rightarrow\left(\frac{u_{m}}{v_{m}}-1\right)t_{m}$ $=\frac{w_{m}}{v_{m}}.$ (95)
@@
Moreover,
@@
As $m$ increases, (101) becomes smaller due to a larger fraction being subtracted from $1$. Thus, $t_{m}$ is decreasing in $m$. Now, we prove the lemma for the base case and apply induction. Base case: For $m=1$, we have
@@
Induction step: Let us assume $q_{m}>t_{m}$ for any $m$. We have
@@
$=\frac{u_{m+1}}{v_{m+1}}(t_{m}-t_{m+1})+t_{m+1}$ (109)
$>t_{m+1}.$ (110)
@@
Thus, for $m=1,2,\dots,\left\lfloor\frac{1}{e^{\epsilon}-1}\right\rfloor$, $q_{m}$ satisfies (93). ∎
Lemma 4
@@
For $m\in\left\{0,1,\dots,\left\lfloor\frac{n}{k}\right\rfloor-1\right\}$, the function $q_{m}$ satisfies the inequality condition
@@
Here, in (118), we substitute (64). (122) follows from the fact that $q_{m}$ is non-increasing and $\frac{1}{k}-\frac{m+1}{n}$ is non-negative for $m=1,2,\dots,\left\lfloor\frac{n}{k}\right\rfloor-1$. Finally, for $q_{m+1}=0$, we have

### v2

@@ B-A Auxiliary Lemmas
Recall the inequality condition in (68),
$(u_{m}-v_{m})q_{m}\leq w_{m},$ (79)
for all $m$. Since $u_{m}-v_{m}<0$ for all $m$, the inequality in 79 is equivalent to
$q_{m}\geq\frac{w_{m}}{u_{m}-v_{m}}.$ (80)
Now, this condition is valid when the right side is positive. Since the denominator is always negative, this happens only when the numerator is also negative, i.e.,
$w_{m}=-\frac{1}{n}-\frac{m}{n}+\frac{m}{n}e^{\epsilon}<0,$ (81)
which happens whenever
$m<\frac{1}{e^{\epsilon}-1}.$ (82)
Let
$t_{m}=\frac{w_{m}}{u_{m}-v_{m}}.$ (83)
Dividing the numerator and denominator by $v_{m}$ and rearranging, we have
$\left(\frac{u_{m}}{v_{m}}-1\right)t_{m}=\frac{w_{m}}{v_{m}}.$ (84)
@@
Moreover, substituting the expressions for $u_{m},v_{m}$ and $w_{m}$ into (83), we have
@@
As $m$ increases, (90) becomes smaller due to a larger fraction being subtracted from $1$. Thus, $t_{m}$ is decreasing in $m$. Now, we prove that $q_{m}\geq t_{m}$ for the base case and apply induction.
Base case: For $m=1$, we have
@@
Induction step: Assume $q_{m}>t_{m}$ for any $m$. Then we have
@@
Rearranging the right side of (97), we have
$q_{m+1}$ $>\frac{u_{m+1}}{v_{m+1}}(t_{m}-t_{m+1})+t_{m+1}.$ (98)
@@
Since $t_{m}$ is strictly decreasing in $m$ and the coefficient of $t_{m}-t_{m+1}$ is positive, $q_{m+1}>t_{m+1}$. Thus, for $m\in\left\{0,1,2,\dots,\left\lfloor\frac{1}{e^{\epsilon}-1}\right\rfloor\right\}$, $q_{m}$ satisfies (78). ∎
Lemma 3.
For $\epsilon>0,k\geq 2,n>k$, and $m<\frac{n}{k}$
$f_{1}(m,q_{m-1})\leq q_{m-1}.$ (99)
Proof:
We may write
$f_{1}(m,q_{m-1})-q_{m-1}$ $=\frac{u_{m}}{v_{m}}q_{m-1}-\frac{w_{m}}{v_{m}}-q_{m-1}$ (100)
$=\left(\frac{u_{m}}{v_{m}}-1\right)q_{m-1}-\frac{w_{m}}{v_{m}}.$ (101)
Thus, to prove that $f_{1}(m,q_{m-1})\leq q_{m-1}$, it suffices to show that the right-hand side of (101) is non-positive. Substituting (65) and (66) into the coefficient of $q_{m-1}$ in (101), we have
$<0,$ (105)
for all $m<\frac{n}{k}$. Similarly, substituting (67) and (66) into the second term in (101), we have
$>0,$ (108)
for $\left\lfloor\frac{1}{e^{\epsilon}-1}\right\rfloor<m<\frac{n}{k}$. Since $q_{m}$ has a lower bound 0 in this range, the right side of (101) is
$\left(\frac{u_{m}}{v_{m}}-1\right)q_{m-1}-\frac{w_{m}}{v_{m}}<0.$ (109)
It still remains to show that the right side is negative for $m<\left\lfloor\frac{1}{e^{\epsilon}-1}\right\rfloor$. We observe from Lemma 2 that in this range of $m$, $q_{m}$ has a non-zero and positive lower bound $t_{m}$ given by the right-hand side of (78). Substituting this bound to the right side of (101), we have
$\left(\frac{u_{m}}{v_{m}}-1\right)q_{m-1}-\frac{w_{m}}{v_{m}}$ $\leq\left(\frac{u_{m}}{v_{m}}-1\right)t_{m-1}-\frac{w_{m}}{v_{m}}.$ (110)
Since $\frac{u_{m}}{v_{m}}-1<0$ and $t_{m}$ is strictly decreasing in $m$, we can write
$\left(\frac{u_{m}}{v_{m}}-1\right)t_{m-1}-\frac{w_{m}}{v_{m}}$ $\leq\left(\frac{u_{m}}{v_{m}}-1\right)t_{m}-\frac{w_{m}}{v_{m}}$ (111)
$=0,$ (112)
where (112) follows from (84). Thus, for all $m<\frac{n}{k}$,
$q_{m}-q_{m-1}$ $=\left(\frac{u_{m}}{v_{m}}-1\right)q_{m-1}-\frac{w_{m}}{v_{m}}\leq 0.$ (113)
Therefore, $f_{1}(m,q_{m-1})\leq q_{m-1}$. ∎
Figure 3: The three cases in the proof of Lemma 4.
Lemma 4.
For $\epsilon>0,k\geq 2,\text{ and }n>k$, for all $m$,
$f_{2}(q_{m-1})\leq q_{m-1}.$ (114)
Proof:
For some fixed $n$, $k$ and $\epsilon$, the function
$f_{2}(q)=\frac{v^{\prime}}{u^{\prime}}q+\frac{w^{\prime}}{u^{\prime}}$ (115)
is linear in $q$, with slope $\frac{v^{\prime}}{u^{\prime}}$ and intercept $\frac{w^{\prime}}{u^{\prime}}$. The slope
$\frac{v^{\prime}}{u^{\prime}}=\frac{e^{\epsilon}\left(\frac{1}{k}-1\right)}{-1+\frac{1}{k}-\frac{1}{n}}=\frac{e^{\epsilon}\left(\frac{1}{k}-1\right)}{\left(\frac{1}{k}-1\right)-\frac{1}{n}}$ (116)
is positive and less than $1$ for $0<\epsilon<\log\left(1+\frac{1}{n-\frac{n}{k}}\right)$. For $\epsilon>\log\left(1+\frac{1}{n-\frac{n}{k}}\right)$, the slope is positive and greater than $1$. The intercept
$\frac{w^{\prime}}{u^{\prime}}=\frac{e^{\epsilon}-1-\frac{1}{n}}{-1+\frac{1}{k}-\frac{1}{n}}$ (117)
is positive for $\epsilon<\log(1+\frac{1}{n})$ and negative for $\epsilon>\log(1+\frac{1}{n})$. Thus, we can consider three regions of $\epsilon$ values as illustrated in Figure 3. Let $q^{*}$ denote the fixed point of $f_{2}$, i.e., the value of $q$ where $f_{2}(q)=q$. Thus
$q^{*}$ $=\frac{\frac{w^{\prime}}{u^{\prime}}}{1-\frac{v^{\prime}}{u^{\prime}}}$ (118)
$=\frac{w^{\prime}}{u^{\prime}-v^{\prime}}.$ (119)
Substituting (62)–(64) into (119) and simplifying, we have
$q^{*}$ $=\frac{e^{\epsilon}-1-\frac{1}{n}}{-1+\frac{1}{k}-\frac{1}{n}-e^{\epsilon}\left(\frac{1}{k}-1\right)}$ (120)
$=\frac{ne^{\epsilon}-n-1}{ne^{\epsilon}-n-1-\frac{n}{k}(e^{\epsilon}-1)}$ (121)
$=\frac{1}{1+\frac{\frac{n}{k}(e^{\epsilon}-1)}{1-n(e^{\epsilon}-1)}}.$ (122)
Recall that $q_{0}=\frac{1}{1+\frac{n}{k}(e^{\epsilon}-1)}$. We now consider each range of $\epsilon$ values and prove that $f_{2}(q_{m-1})\leq q_{m-1}$ holds for each case.
Case I: $0<\epsilon<\log(1+\frac{1}{n})$. In this region, we have $0<\frac{v^{\prime}}{u^{\prime}}<1$ and a positive intercept $\frac{w^{\prime}}{u^{\prime}}$. Thus, $q^{*}$ is positive. Moreover, $f_{2}(q)\leq q$ for $q\geq q^{*}$, so to prove the lemma it is enough to show that $q_{m}\geq q^{*}$ for all $m$. We will prove this by induction. For the base case, consider $m=0$. From (122), we observe that $0<1-n(e^{\epsilon}-1)<1$ in this range of $\epsilon$, which implies $q^{*}<q_{0}$. This proves the base case. Now, for the inductive step assume that $q_{m-1}\geq q^{*}$. Thus we have
$q_{m}=\max\{0,f_{1}(m,q_{m-1}),f_{2}(q_{m-1})\}\geq f_{2}(q_{m-1})\geq q^{*}.$ (123)
This completes the proof for this case.
Case II: $\log(1+\frac{1}{n})\leq\epsilon\leq\log\left(1+\frac{1}{n-\frac{n}{k}}\right)$. In this region, the slope satisfies $0<\frac{v^{\prime}}{u^{\prime}}\leq 1$, but the intercept satisfies $\frac{w^{\prime}}{u^{\prime}}\leq 0$. Consequently, $f_{2}(q)\leq q$ for all $q\geq 0$, which proves the lemma in this case since by construction $q_{m-1}\geq 0$ for all $m$.
Case III: $\epsilon>\log\left(1+\frac{1}{n-\frac{n}{k}}\right)$. In this region, we have $\frac{v^{\prime}}{u^{\prime}}>1$ and a negative intercept $\frac{w^{\prime}}{u^{\prime}}$. Thus, $q^{*}$ is positive. Since $f_{2}(q)\leq q$ for $q\leq q^{*}$, it is enough to show that $q_{m}\leq q^{*}$ for all $m$. We prove this by induction. For the base case, consider $m=0$. We observe that $1-n(e^{\epsilon}-1)<0$ in this range of $\epsilon$. This implies that
$1+\frac{n}{k}(e^{\epsilon}-1)$ $>1+\frac{\frac{n}{k}(e^{\epsilon}-1)}{1-n(e^{\epsilon}-1)}.$ (124)
Therefore, $q_{0}<q^{*}$. Now, for the inductive step, assume for all $m^{\prime}<m$, $q_{m^{\prime}}\leq q^{*}$. Thus, $f_{2}(q_{m^{\prime}})\leq q_{m^{\prime}}$ for all $m^{\prime}<m$. From Lemma 3, we also know that $f_{1}(m^{\prime},q_{m^{\prime}-1})\leq q_{m^{\prime}-1}$ for all $m^{\prime}$. Thus, for all $m^{\prime}\leq m$,
$q_{m^{\prime}}=\max\{0,f_{1}(m^{\prime},q_{m^{\prime}-1}),f_{2}(q_{m^{\prime}-1})\}\leq\max\{0,q_{m^{\prime}-1}\}=q_{m^{\prime}-1}.$ (125)
That is, $q_{m^{\prime}}$ is non-increasing in $m^{\prime}$ for all $m^{\prime}\leq m$, which means $q_{m}\leq q_{0}$. As we have already shown that $q_{0}\leq q^{*}$, this proves that $q_{m}\leq q^{*}$, which completes the proof of the lemma. ∎
Lemma 5.
@@
For $\epsilon>0,k\geq 2,n>k$, $m\in\left\{0,1,\dots,\left\lfloor\frac{n}{k}\right\rfloor-1\right\}$, and $u_{m},v_{m}$ and $w_{m}$ as defined in (65)–(67), the function $q_{m}$ satisfies the inequality condition
@@
Here, (137) follows from the fact that $q_{m}$ is non-increasing and $\frac{1}{k}-\frac{m+1}{n}$ is non-negative for $m=1,2,\dots,\left\lfloor\frac{n}{k}\right\rfloor-1$. Finally, for $q_{m+1}=0$, we have